OpenStudy (anonymous):

Simplify the following expression: 5x − 6y + 5z − x − 12y + 2z 6x − 18y + 7z 4x − 18y + 7z 4x − 6y + 7z 6x − 6y + 7z

OpenStudy (anonymous):

@imqwerty

Nnesha (nnesha):

combine like terms if the variables and exponent are the same then you can add/subtract their coefficients

Nnesha (nnesha):

here is an example 6x+4x variables are same with same exponent so you can add 6 and 4 `6x+4x` = (6+4)x=10x

Nnesha (nnesha):

\[5x − 6y + 5z − x − 12y + 2z \] what are like terms in this expression ?

imqwerty (imqwerty):

1 Gather like terms (5x−x)+(−6y−12y)+(5z+2z) 2 Simplify 4x−18y+7z

Nnesha (nnesha):

you're not new here but again would be great of u at least give them a chance to try it first. they can do it! they are smart!

OpenStudy (anonymous):

@imqwerty what is it?

OpenStudy (anonymous):

nvm i see it .. thank u

imqwerty (imqwerty):

(: yw